OpenStudy (anonymous):

If f(x) = 3x2 – 2x + 10 and g(x) = 8x – 5, what is f(x) + g(x)?

OpenStudy (anonymous):

3x2 – 2x + 10 +8x – 5 =3x2 +6x +5

OpenStudy (anonymous):

3x2 +6x +5

mathslover (mathslover):

f(x) = \(3x^2-2x+10\) g(x) = \(8x-5\) f(x)+g(x) = \(\huge{3x^2-2x+10+8x-5}\) can u solve this now ?

OpenStudy (theviper):

\[\Huge{\color{blue}{3x^2+6x+5}}\]

OpenStudy (anonymous):

consider f(x) + g(x) 3x^2-2x +10 +8x -5 3x^2 +6x + 5

OpenStudy (anonymous):

hey guys can you find the value of A in tan(A -45) =1/3

OpenStudy (theviper):

Maybe @mathslover can find:)

OpenStudy (anonymous):

hahaha... why not you

OpenStudy (theviper):

I didn't learn Trigonometry:(

OpenStudy (anonymous):

ok am good at trigonometric will you need a help